Show that the X solutions of the simultaneous equations
Y = X - 2
xy = 3
Are solutions of the equation
X(squared) - 2x - 3 = 0

Answer by jim_thompson5910(35256)

Note: to say "x squared" we type "x^2"


x*y = 3 ... start with the second equation


x*(x-2) = 3 ... replace every y with (x-2). This is valid because the first equation is y = x-2


x*x + x*(-2) = 3 ... distribute


x^2 - 2x = 3 ... multiply


x^2 - 2x - 3 = 0 ... subtract 3 from both sides